### v1

# Symplectic mapping class relations from pencil pairs

Abstract.
We describe symplectic mapping class relations between products of positive Dehn twists along Lagrangian spheres in Weinstein $4$-manifolds, all of which are affine $\mathbb{C}$ varieties. The relations are obtained by applying classification results for Fano $3$-folds and polarized $K_{3}$ surfaces of small genus to a general methodology – finding pencil pairs.
1. Introduction
This article describes a general method for finding symplectic mapping class relations between products of positive Dehn twists on (affine algebraic) Weinstein manifolds $F$ of $\dim=2n$ by studying complex projective varieties of $\dim_{\mathbb{C}}=n+1$. The general method consists of finding pencil pairs and is known to experts. It recovers known mapping class relations on punctured Riemann surfaces such as the lantern relation [AS08, J79], described in §3.3 and depicted in Figure 1. In formally defining pencil pairs and working out some basic techniques for finding them (see §4), we hope to inspire specialists in symplectic topology and algebraic geometry to further study the case $\dim F\geq 4$, for which few examples of such symplectic mapping class relations are known. Therefore our exposition will cover some relevant basics from both subject areas. Our search for mapping class relations on high-dimensional Weinstein manifolds is motivated by recent advances in high-dimensional contact topology. See §2.3.
[…]
Figure 1. The lantern relation states that a product of Dehn twists along circles shown on the left is isotopic (relative to the boundary of the $4$-holed sphere) to a product of Dehn twists along the circles shown on the right.
In [O21] Oba finds a symplectic mapping class relation on a Weinstein $4$-manifold by applying the classification of del Pezzo surfaces [F80] to find a pencil pair. Our main application applies classification results for polarized $K_{3}$ varieties [M87, SD74] and Fano $3$-folds [I77, I77, MM82] to find new pencil pairs, and so new examples of mapping class relations on Weinstein $4$-manifolds. Here is one such family of examples.
Theorem 1.0.1.
Let $Z\subset\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{2}$ be the $K_{3}$ surface cut out by a smooth complete intersection of $\deg=(1,1,2)$ and $\deg=(1,1,1)$ divisors. Let $(F,\beta_{F})$ be the Weinstein manifold given by the complement of a $\deg=(1,1,1)$ divisor in $Z$, with the symplectic structure determined by a $\deg=(1,1,1)$ line bundle. Then there is a compactly supported symplectomorphism $\tau_{\partial F}\in\pi_{0}\operatorname{Symp}^{c}(F,d\beta_{F})$ which can be factored into a product of $m$ Dehn twists along Lagrangian spheres in $(F,\beta_{F})$, where
$m=18,32,34,36,40.$
We found six such families of mapping class relations in total, each giving rise to an infinite family of mapping class relations by a generally-applicable “cabling” argument described in Lemma 4.2.1. One of these families (see §5.3) recovers the relation of [O18]. Because the number of Dehn twists in our relations are all distinct, they cannot be derived from well-known braid relations [KS02]. See §2.2.
1.1. Conventions
All varieties in this article are smooth and defined over $\mathbb{C}$. Throughout $\dim$ and $\dim_{\mathbb{C}}$ refer to real and complex dimension. On a closed manifold $X$ we use $\operatorname{PD}_{X}$ for Poincaré duality in either direction $H_{k}(X)\leftrightarrow H^{\dim X-k}(X)$. All of our symplectic manifolds will be complex projective varieties $X,Z,B,\mathcal{M}\subset\mathbb{P}^{n}$ or affine varieties $F,W\subset\mathbb{C}^{n}$, with their symplectic structures inherited from the ambient space.
1.2. Acknowledgements
We thank Eduardo Alves da Silva, Yankı Lekili, and Takahiro Oba for interesting discussions. We’re especially grateful to Marco Golla for discussions around complex surfaces, reviewing an early draft of the article, and for pointing out the $l>1$ examples of §3.3. We thank Cofund MathInGreaterParis and the Fondation Mathématique Jacques Hadamard for supporting us as a member of the Laboratoire de Mathématiques d’Orsay at Université Paris-Saclay. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034255.
2. Background
2.1. Symplectic mapping class relations
Let $(F,\beta_{F})$ be an exact symplectic manifold, meaning that $\beta_{F}\in\Omega^{1}(F)$ is such that $\omega_{F}=d\beta_{F}$ is symplectic. We require that $F$ is a compact Liouville domain – meaning that the vector field $X_{F}$ determined by $d\beta_{F}(X_{F},\ast)=\beta_{F}$ points transversely out of $\partial F$ – or the completion of a compact Liouville domain, meaning in any case that it has finite type. We say that $(F,\beta_{F})$ is Weinstein if $X_{F}$ is gradient-like for some function $f$ on $F$. See [CE12] for details and background. The canonical examples of Weinstein manifolds are affine algebraic varieties $F\subset\mathbb{C}^{N}$ with $f=\left\lVert z\right\rVert^{2}$ and $\beta=-df\circ J_{0}$, where $J_{0}$ is the standard complex structure on $\mathbb{C}^{N}$. The group of compactly supported symplectomorphisms will be denoted $\operatorname{Symp}^{c}(F,\omega_{F})$ and we are interested in studying $\pi_{0}\operatorname{Symp}^{c}(F,\omega_{F})$.
An exact Lagrangian sphere $L\subset F$, where $\dim F=2n$ is a sphere $\mathbb{S}^{n}$ for which $\beta_{F}|TL=df$ for some $f\in\mathcal{C}^{\infty}(L)$. A framing on such a sphere is an identification of the sphere with the boundary of a disk, up to homotopy. We give a quick review of Dehn twists about such $L$, which appear in nature as monodromies of Milnor fibers of the singularities $\sum_{0}^{n}z_{j}^{2}$ on $\mathbb{C}^{n+1}$ [Arn95]. According to the Weinstein neighborhood theorem, every such $L$ has a neighborhood of the form $(N_{F}(L),\beta_{F})=(\mathbb{D}^{\ast}\mathbb{S}^{n},pdq+df)\implies d\beta_{F}|_{N_{F}(L)}=dp\wedge dq$ where $pdq$ is the canonical $1$-form on a cotangent disk bundle $\mathbb{D}^{\ast}\mathbb{S}^{n}$ of the sphere and $f\in\mathcal{C}^{\infty}(\mathbb{D}^{\ast}\mathbb{S}^{n})$ is some smooth function. In particular, writing $\bullet$ for the intersection pairing on $H_{n}(L)$, $[L]\bullet[L]$ is 00 for $n$ odd. When $n$ is even, $[L]\bullet[L]=-2$ meaning that $[L]$ must be primitive in $H_{n}(F)$. Let’s assume the disk bundle is defined using the standard round metric of radius $1$. Then can define a Dehn twist $\tau_{L}$ locally as $\tau_{L}=(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}\operatorname{Flow}^{1}_{X_{b}}$ where $b=b(\left\lVert p\right\rVert)$ is a function which equals 00 near $\left\lVert p\right\rVert=0$ and $\pi$ near $\left\lVert p\right\rVert=1$, $X_{b}$ is its Hamiltonian vector field $db=dp\wedge dq(\ast,X_{b})$, and $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ is the symplectomorphism induced by the antipodal map $q\mapsto-q$ on $\mathbb{S}^{n}$. In words, $\operatorname{Flow}^{1}_{X_{b}}$ applies the time $b(\left\lVert p\right\rVert)$ geodesic flow to (co)vectors of length $\left\lVert p\right\rVert$. In particular, when $\left\lVert p\right\rVert=1$, the time $1$-flow will send a vector over $q\in\mathbb{S}^{n}$ to a vector over $-q$ and the $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ correction ensures that $\tau_{L}$ is the identity near $\partial\mathbb{D}^{\ast}\mathbb{S}^{n}$. Extend $\tau_{L}$ to $F\setminus N_{F}(L)$ as the identity. The dependence on $b$ is irrelevant when we view $\tau_{L}\in\pi_{0}\operatorname{Symp}^{c}$. It can be seen by drawing pictures that when $\dim F=2$, $\tau_{L}$ coincides with the usual notion of a Dehn twist along a simple closed curve on a Riemann surface. The action of $\tau_{L}$ on $H_{\ast}(F)$ is given by the Picard-Lefschetz formula
(2.1.1) $(\tau_{L})h=\begin{cases}h&h\in H_{\neq n}(F),\\ h+(-1)^{\frac{1}{2}(n+1)(n+2)}(h\bullet[L])[L]&h\in H_{n}(F).\end{cases}$
In the cases $n=1,3$, $\tau_{L}^{2}$ is smoothly – but not symplectically – isotopic to the identity [Av21, Se99]. See [KRW23] for more on powers of Dehn twists.
A positive factorization of $\phi\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{F})$ is a tuple $\vec{L}=(L_{1},\cdots L_{m})$ of framed, exact Lagrangian spheres for which
$\phi=\tau_{L_{m}}\cdots\tau_{L_{1}}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{F}).$
We say that $m$ is the length of the positive factoriation, written
$|\vec{L}|=m.$
Two positive factorizations $\vec{L}$ and $\vec{L}^{\prime}=(L_{1}^{\prime},\cdots L_{m^{\prime}}^{\prime})$ of a $\phi$ are deformation equivalent if $m=m^{\prime}$ and for each $i$, $L_{i}$ is Hamiltonian isotopic to $L_{i}^{\prime}$, whence we write $\vec{L}=\vec{L}^{\prime}$. We are interested in such deformation equivalence classes of $\vec{L}$ as the provide instructions for building Weinstein manifolds of $\dim=\dim F+2$ as described in §2.3.
Definition 2.1.1.
A positive relation on $F$ is a pair of positive factorizations $\vec{L},\vec{L}^{\prime}$ of some $\phi$.
2.2. Known relations and counting twists
There are a few well-known ways to modify a positive factorization – and so determine positive relations:
  1. (1)
    The disjoint union relation states that $L\cap L^{\prime}=\emptyset\implies\tau_{L}\tau_{L^{\prime}}=\tau_{L^{\prime}}\tau_{L}$.
  2. (2)
    The conjugacy relation states that for any two $L$ and $L^{\prime}$, $\tau_{L}\tau_{L^{\prime}}=\tau_{\tau_{L}L^{\prime}}\tau_{L}$.
  3. (3)
    The braid relation states that if $L$ and $L^{\prime}$ intersect transversely in a single point, $\tau_{L}\tau_{L^{\prime}}\tau_{L}=\tau_{L^{\prime}}\tau_{L}\tau_{L^{\prime}}$.
The first two are clear. The last can be verified directly by studying Lagrangian spheres in $A_{2}$ Milnor fibers and gives rise to an embedding of the braid group on $m$ strands into the $\pi_{0}\operatorname{Symp}^{c}$ of an $A_{m}$ Milnor fiber [KS02]. When $\dim F=4$, this embedding is an isomorphism [Ev11, W14].
We seek to single out the positive factorizations which are not derived from braid groups and so are “not obvious”. Let $T=T(F,\beta_{F})$ be the free group on Hamiltonian isotopy classes of exact framed Lagrangian spheres. Consider the normal subgroup $B=B(F,\beta_{F})\subset T$ generated by $[L][L^{\prime}][L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ which are disjoint, $[L][L^{\prime}][L]^{-1}[\tau_{L}L^{\prime}]^{-1}$ for any $L,L^{\prime}$, and $[L][L^{\prime}][L][L^{\prime}]^{-1}[L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ intersecting transversely in $F$ a single point. We have canonical homomorphisms
$\tau:T\rightarrow\pi_{0}\operatorname{Symp}(F,\omega_{F}),\quad\tau[L]=\tau_{L},\quad p:T\rightarrow\mathbb{Z},\quad p[L]=1$
which factor through $T/B$. In words, $\tau$ sends formal products of exact Lagrangian spheres (and their inverses) to their products of their positive (and negative) Dehn twists and $p$ measures the sum of the exponents. So for a positive factorization $\vec{L}$, $p\vec{V}=|\vec{L}|$. By construction $\ker\tau$ gives all relations in the subgroup $\operatorname{im}\tau$ of compactly supported symplectomorphisms generated by Dehn twists.
Each positive factorization $\vec{L}$ of some $\phi$ determines elements of $T$ and $T/B$ by $[\vec{L}]=[L_{k}]\cdots[L_{1}]$. A positive relation $\vec{L},\vec{L}^{\prime}$ is non-braid-like if $[\vec{L}]\neq[\vec{L}^{\prime}]\in T/B$. It’s clear that if $|\vec{L}|\neq|\vec{L}^{\prime}|$, then the relation is non-braid-like. So examples of non-braid-like relations include the lantern relations and chain relations on punctured Riemann surfaces, cf. [OS04b, §15].
We are interested in finding non-braid like positive relations in the case $\dim F>2$. In a way which is made precise in [K22], the existence of such relations for $\dim F>2$ is more restrictive than the $\dim F=2$ case. Here is an easy way to see that relations in the $\dim F=2$ case (such as the lantern, which has $p=\pm 1$) cannot be straightforwardly generalized to higher dimensions.
Lemma 2.2.1.
Suppose that $4$ divides $\dim F$ and $\vec{L}\in T$ is such that $\tau\vec{L}$ acts trivially on $H_{n}(F)$ (which of course must be satisfied if $\tau\vec{L}=\operatorname{Id}_{F}$). Then $p\vec{L}\in 2\mathbb{Z}$.
Proof.
Consider the action of a single $\tau_{L}$ on the determine line $\bigwedge^{top}H_{n}(F)$. Since $[L]$ is primitive, we can extend $[L]$ to a basis of $H_{n}(F)$ so that $\tau_{L}\in\operatorname{Aut}(H_{n})$ can be expressed as a matrix. The first row has $-1$ in the first entry (combining $[L]\bullet[L]=-\chi(\mathbb{S}^{n})=-2$ with Equation (2.1.1)) with the rest of the entries being zeros. In the $j$th row for $j>1$, Equation (2.1.1) determines the first entry, the $j$th entry is $1$, and all other entries are zero. So $\tau_{L}$ acts on the determinant line by $-1$. For a general $\vec{L}\in T$, $\tau\vec{L}$ acts on the determinant line by $(-1)^{p\vec{L}}$. ∎
2.3. Some contact-topological motivations
We outline the role of positive factorizations and positive relations in modern contact topology. This is an advertisement aiming at those readers who already have some exposure to contact topology and so the exposition will be light on details.
Given a positive factorization $\vec{L}$ as above we can build a Liouville manifold $W=W_{\vec{L}}$ of $\dim W_{\vec{L}}=\dim F+2$ as follows: Take $W_{0}$ to be $(\mathbb{D}\times F,r^{2}d\theta+\beta_{F})$ with the corner $\partial\mathbb{D}\times\partial F$ rounded. The $\Lambda_{j}=(e^{j2\pi i/m},L_{j})$ will be framed Legendrian spheres in the contact boundary of $W_{0}$ (after a slight perturbation of the boundary of $W_{0}$ to manage the exact $1$-forms $\beta_{F}|_{TL_{i}}$). Then $W$ is obtained by attaching $\operatorname{ind}=\frac{1}{2}\dim W_{0}$ Weinstein handles along the $\Lambda_{j}$. It follows that
$\chi(W)=\chi(F)+(-1)^{\frac{1}{2}\dim W}|\vec{L}|.$
When $(F,\beta_{F})$ is Weinstein, it follows that $(W,\beta_{W})$ is as well. In any case, the boundary of $W$ is naturally a contact manifold $(\Gamma,\xi_{\Gamma})$ described as an open book $\mathcal{OB}(F,\beta_{F},\phi)$ with page $(F,\beta_{F})$ and monodromy $\phi$, whose contact structure depends only on $\phi\in\pi_{0}\operatorname{Symp}^{c}(F,d\beta_{F})$ [BHH23, Gi02, TW75]. It’s known that any Weinstein manifold of $\dim\geq 4$ can be constructed in this way with $(F,\beta_{F})$ Weinstein [BHH23, GP17].
It follows that a positive relation $\vec{L},\vec{L}^{\prime}$ gives us two possibly distinct Liouville fillings $(W=W_{\vec{L}},\beta)$ and $(W^{\prime}=W_{\vec{L}^{\prime}},\beta^{\prime})$ of the same $(\Gamma,\xi_{\Gamma^{\prime}})$. So positive relations help us to construct distinct Liouville and Weinstein fillings of a given contact manifold, cf. [O18, OS04a].
Provided a positive relation $\vec{L},\vec{L}^{\prime}$, the presentation of the contact boundaries of $W$ and $W^{\prime}$ as open books with the same monodromy give us an identifications of these boundaries. As described in [Br23], we can use these fillings to determine a $\tau$-invariant contact structure $\xi$ on $N=(-\epsilon,\epsilon)_{\tau}\times S,S=W\cap_{\Gamma}W^{\prime}$. We say that $S$ is a convex hypersurface with standard neighborhood $(N,\xi)$. Identifying $S=\{0\}\times S$, the Euler class of $\xi|_{S}$ is computed
(2.3.1) $\int_{S}e(\xi)=\chi(W)-\chi(W^{\prime})=(-1)^{\frac{1}{2}\dim W}\left(|\vec{L}|-|\vec{L}^{\prime}|\right).$
Combining [HH19] and [Br23], any oriented hypersurface in a contact manifold $(M,\xi)$ of $\dim M\geq 5$ has such a neighborhood (determined by a positive relation on a Weinstein $(F,\beta_{F})$) after applying a $\mathcal{C}^{0}$ small perturbation to the hypersurface. Since convex hypersurfaces exist in abundance by [HH19], so should positive relations. Moreover, Equation (2.3.1) tells us that there are lots of positive relations $\vec{L},\vec{L}^{\prime}$ which are not braid like. Here are some examples.
Example 2.3.1.
For a fixed $n\in\mathbb{Z}_{>2}$ and a partition $\vec{n}=(n_{1},\cdots,n_{m}),n_{i}\leq n_{i+1},\sum n_{i}=n$, let $\mathbb{P}^{\vec{n}}=\mathbb{P}^{n_{1}}\times\cdots\times\mathbb{P}^{n_{m}}$. Then $T\mathbb{P}^{\vec{n}}$ defines an almost contact structure on the $\dim=2n+1$ manifold $M^{\vec{n}}=\mathbb{S}^{1}\times\mathbb{P}^{\vec{n}}$. By [BEM15] – which includes a definition of almost contact structures – there is a (genuine) contact structure $\xi^{\vec{n}}$ on $M^{\vec{n}}$ which is homotopic to $T\mathbb{P}^{\vec{n}}\subset TM^{\vec{n}}$ through almost contact structures.
Let $(M,\xi)$ be the contact connected sum of the $(M^{\vec{n}},\xi^{\vec{n}})$, over all partitions $\vec{n}$, the number of which we’ll write as $N_{n}$. For $\vec{r}=(r_{\vec{n}})\in\mathbb{Z}^{N_{n}}$ consider the homology class $h_{\vec{r}}\in H_{2n}(M)$ given by the sum of the $r_{\vec{n}}[\{\theta\}\times\mathbb{P}^{\vec{n}}]$, with each summand living in the image of $H_{2n}(M^{\vec{n}})\simeq H_{2n}(M^{\vec{n}}\setminus\mathbb{D}^{2n+1})\rightarrow H_{2n}(M)$. It’s not difficult to see that $h_{\vec{r}}$ has a connected, embedded, oriented representative, $S_{0}$. By [HH19] we can apply a $\mathcal{C}^{0}$-small perturbation to $S_{0}$ to obtain a $S\subset M$ which is convex with neighborhood $(N,\xi)$. The contact structure $\xi|_{S}$ gives $S$ the structure of a stably almost complex manifold whose cobordism class is $[S]=\sum r_{\vec{n}}\mathbb{P}^{\vec{n}}\in\Omega^{U}_{2n}$. Since the $\mathbb{P}^{\vec{n}}$ additively generate $\Omega^{U}_{2n}$ (cf. [MiS74]), we see that any almost complex cobordism class can be represented by a convex hypersurface.
By [Br23] there is some Weinstein $(F,\beta_{F})$ of $\dim=2n-2$ having a positive relation $\vec{L},\vec{L}^{\prime}$ determining $(N,\xi)$. Computing $(-1)^{n}\int_{S}e(\xi)=(-1)^{n}\int_{S}c_{n}(\xi)$ gives us the difference in the lengths of the positive relation. Note that the image of $\int_{\ast}c_{n}:\Omega^{U}_{2n}\rightarrow\mathbb{Z}$ is $2\mathbb{Z}$ when $n$ is odd and $\mathbb{Z}$ when $n$ is even. Taking $m=n-1$, we see that for every $m\geq 2$ there are Weinstein $(F,\beta_{F})$ of $\dim=2m$ having positive relations with $|\vec{L}|-|\vec{L}^{\prime}|$ being any integer when $m$ is odd and any even integer when $m$ is even.
Unfortunately it is very difficult to explicitly describe such $(N,\xi)$, even for convex hypersurfaces in well-studied $(M,\xi)$. In searching for symplectic mapping class relations, we seek to have some interesting examples of convex hypersurfaces at least for which the $(F,\beta_{F})$ can be explicitly described (eg. as affine varieties). In a follow-up article, we will study the $(N,\xi)$ determined by some of the positive relations in this article using the techniques of [Av23a].
2.4. Review of algebraic Lefschetz pencils
Let $\operatorname{\mathcal{L}}$ be an ample line bundle on a closed variety $X$. The pair $(X,\operatorname{\mathcal{L}})$ – a polarization of $X$ – determines an immersion $\phi_{\operatorname{\mathcal{L}}}:X\rightarrow\mathbb{P}^{m}$, $m=\dim_{\mathbb{C}}H^{0}(X,\operatorname{\mathcal{L}})-1$ such that $\phi_{\operatorname{\mathcal{L}}}^{\ast}\mathcal{O}_{1}=\operatorname{\mathcal{L}}$ and for which $\omega_{\operatorname{\mathcal{L}}}=\phi_{\operatorname{\mathcal{L}}}^{\ast}\omega_{\mathbb{P}^{m}}$ is symplectic. Here $\omega_{\mathbb{P}^{m}}$ is the Fubini-Study form, normalized to have volume $1$. When $\phi_{\operatorname{\mathcal{L}}}$ is an embedding, $\operatorname{\mathcal{L}}$ is very ample. Replacing $\operatorname{\mathcal{L}}$ with $\operatorname{\mathcal{L}}^{\otimes k}$ for $k\in\mathbb{Z}_{>0}$ gives more ample line bundles.
A pair $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ of distinct holomorphic sections of $\operatorname{\mathcal{L}}$ yields
  1. (1)
    subvarieties $Z_{i}=\{\boldsymbol{\mathfrak{s}}_{i}=0\}\subset X$ (ie. an intersection of $X\subset\mathbb{P}^{m}$ with a hyperplane determined by a section of $\mathcal{O}_{1}\rightarrow\mathbb{P}^{m}$) with $\operatorname{expdim}_{\mathbb{C}}V_{i}=\dim_{\mathbb{C}}X-1$,
  2. (2)
    a base locus $B=B_{\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1}}=Z_{0}\cap Z_{1}$ with $\operatorname{expdim}_{\mathbb{C}}B=\dim_{\mathbb{C}}X-2$, and
  3. (3)
    a holomorphic function $\pi=\boldsymbol{\mathfrak{s}}_{1}/\boldsymbol{\mathfrak{s}}_{0}:X\setminus Z_{0}\rightarrow\mathbb{C}$.
We say that $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a Lefschetz pencil for $\operatorname{\mathcal{L}}$ if the following criteria are met:
  1. (1)
    The $Z_{i}$ are transversely cut out, so that $\operatorname{PD}_{X}(\omega_{\operatorname{\mathcal{L}}})=[Z_{i}]\in H_{\dim X-2}(X)$,
  2. (2)
    The $Z_{i}$ are transverse to one another, so $B$ is smooth with $\operatorname{PD}_{Z_{i}}(\omega_{\operatorname{\mathcal{L}}})=[B]\in H_{\dim X-4}(Z_{i})$.
  3. (3)
    The function $\pi$ has only non-degenerate critical points $\operatorname{Crit}(\pi)\subset X\setminus Z_{0}$ with pair-wise distinct critical values in $\operatorname{Critv}(\pi)\subset\mathbb{C}$.
It follows from Bertini’s theorem that a generic pair of sections of an ample bundle will determine a pencil. The proof of the following lemma is well-known. We give a quick sketch since variations of the argument will be used throughout the article.
Lemma 2.4.1.
The symplectomorphism type of the $(Z_{1},\omega_{\operatorname{\mathcal{L}}},B)$ associated to a pencil for $\operatorname{\mathcal{L}}$ is independent of sections $\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1}$.
Proof.
Consider a generic $T\in[0,1]$ family $(\boldsymbol{\mathfrak{s}}_{0}^{T},\boldsymbol{\mathfrak{s}}_{1}^{T})$ of pairs of sections of $\operatorname{\mathcal{L}}$. By Bertini’s theorem the set of such pairs for which either $\{\boldsymbol{\mathfrak{s}}_{0}=0\}$ or $\{\boldsymbol{\mathfrak{s}}_{0}=\boldsymbol{\mathfrak{s}}_{1}=0\}$ has complex codimension $\geq 1$ (real codimension $\geq 2$), so we can assume that the path of pairs of sections determines a path of pencils. Applying a Darboux-Moser-Weinstein argument [MS99] to the $T$-family establishes the claim. ∎
When $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a Lefschetz pencil then $\pi$ is a Lefschetz fibration. We will outline the essential properties and point to the references, eg. [AS08, D96b, OS04b, Se08], for a full definition. The complement
$W=X\setminus Z_{0}$
is an affine variety and so is naturally a Weinstein manifold whose underlying symplectic form is $\omega_{\operatorname{\mathcal{L}}}$. So are the regular fibers $F_{z}:=\pi^{-1}(z),z\in\mathbb{C}\setminus\operatorname{Critv}(\pi)$ of $\pi$. The closure of each $F_{z}$ in $X$ intersects $Z_{0}$ along $B$. By considering holonomy along paths in $\mathbb{C}\setminus\operatorname{Critv}(\pi)$, each $F_{z}$ is deformation equivalent to
$F:=F_{0}=Z_{1}\setminus B.$
The pencil $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ determines a compactly supported symplectomorphism of $(F,\omega_{\operatorname{\mathcal{L}}})$ in two ways. First, we study the monodromy of $\pi$. Let $\eta_{i}$ be a collection of embedded, oriented, $[0,1]$-parameterized paths in $\mathbb{C}$ indexed by the $z_{i}\in\operatorname{Critv}(\pi)$, begin at $0\in\mathbb{C}$ and ending at $z_{i}$, and disjoint away from $0\in\mathbb{C}$. We require that the $i$ are ordered so that tangent vectors of the $\eta_{i}$ at 00 are ordered counterclockwise. Each $\eta_{i}$ determines a Lagrangian disk $D_{i}$ in $W$ for which $\pi D_{i}=\eta$ and for which $L_{i}=\partial D_{i}$ is an exact Lagrangian sphere in $F$. The $\eta_{i}$ are vanishing paths and the $D_{i}$ are their associated Lefschetz thimbles which provide framings of the $L_{i}$. Let $\gamma\subset\mathbb{C}\setminus\operatorname{Critv}(\pi)$ be a parameterized simple closed curve based at 00 and oriented counterclockwise, otherwise disjoint from the $\eta_{i}$, encircling all of the $\eta_{i}$, and encircling $\eta_{1}$ first. Then the monodromy morphism $\Phi_{\gamma}\in\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$ can be expressed as a product of Dehn twists $\Phi_{\gamma}=\tau_{L_{\#\operatorname{Crit}(\pi)}}\cdots\tau_{L_{1}}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega)$, generalizing [Arn95]. Modifying the choices of $(\eta_{i})$ corresponds to application of conjugacy relations on the $\vec{L}$.
Second, note that a collar neighborhood of the ideal boundary $\Gamma$ of $F$ is identical to a $\mathbb{D}$-bundle neighborhood of $B\subset Z_{1}$ with its zero section removed. So we can associate to this data a fibered Dehn twist $\tau_{\partial F}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$, cf. [CDvK14, O21]. We give a brief review: Our collar neighborhood of $F$ is given as a symplectization $([0,\infty)_{s}\times\Gamma,e^{s}\alpha)$ determined by a contact form $\alpha$ on $\Gamma$ whose Reeb flow $R_{\alpha}$ is $2\pi$-periodic. Using this collar we can write
$\tau_{\partial F}(s,x)=\left(s,\operatorname{Flow}_{R_{\alpha}}^{g(s)}(x)\right)$
for a smooth function $g=g(s)$ which is increasing with $g(0)=0$ and $g(s)=2\pi$ for $s\geq 1$. Then $\tau_{\partial F}$ extends to the rest of $F$ as the identity map. The induced action on $H_{\ast}(F)$ by $\tau_{\partial F}$ is then clearly the identity. Observe that when $\dim F=2$, $\tau_{\partial F}$ always exists and is a product of positive Dehn twists along simple curves, one parallel to each boundary component of $F$. These simple closed curves are smoothly isotopic to exact Lagrangians iff $F$ has at least two boundary components.
The first statement of following theorem is proved in [Aro01, Go04], with special cases worked out in [AA16, O21]. The count of critical points as a weighted sum of Euler characteristics follows easily from the Lefschetz hyperplane theorem, cf. [O21, Theorem 4.5].
Theorem 2.4.2.
In the above notation
$\tau_{L_{|\vec{L}|}}\cdots\tau_{L_{1}}=\tau_{\partial F}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$
In particular, $\tau_{\partial F_{0}}$ is isotopic to a symplectomorphism determined by a product of
(2.4.1) $|\vec{L}|=\#\operatorname{Crit}\pi=(-1)^{\dim_{\mathbb{C}}X}\left(\chi(X)-2\chi(Z_{1})+\chi(B)\right)$
positive Dehn twists along Lagrangian spheres.
It’s easy to see that not all $\tau_{\partial F}$ can be factorized as products of Dehn twists along exact Lagrangian spheres. For example, take $(F,\beta_{F})$ to be the complement of a $\deg\geq 2$ hypersurface in $\mathbb{P}^{n},n>1$. The ideal boundary of $F$ is a $\mathbb{S}^{1}$-fibered contact manifold so that $\tau_{\partial F}$ is defined. However $F$ contains no Lagrangian spheres, since $\mathbb{P}^{n}$ contains no Lagrangian spheres – a fact which is obvious for homological reasons when $n$ is even and non-trivial when $n$ is odd [Se00]. Addition examples due to Biran and Giroux are described in [CDvK14].
3. Symplectic mapping class relations from pencil pairs
In this section we define pencil pairs and then appeal to Theorem 2.4.2 to show how they determine positive relations. Then we will describe some low-dimensional examples.
3.1. Pencil pairs
Definition 3.1.1.
Let $X^{\pm}$ be closed varieties with ample line bundles $\operatorname{\mathcal{L}}^{\pm}$ equipped with Lefschetz pencils $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$ of $\operatorname{\mathcal{L}}^{\pm}$ and associated subvarieties $B^{\pm}\subset Z^{\pm}_{i}\subset X^{\pm}$. If there is a symplectomorphism
$\phi:(Z_{1}^{+},\omega_{\operatorname{\mathcal{L}}^{+}})\rightarrow(Z_{1}^{-},\omega_{\operatorname{\mathcal{L}}^{-}}),\quad\phi(B^{+})=B^{-},$
then we say that $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
Our notation indicates that the pencil pair depends only on the $\operatorname{\mathcal{L}}^{\pm}$ rather than the specified sections $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$, which follows from Lemma 2.4.1.
A pencil pair together with choices of vanishing paths associated to the $\pi^{\pm}=\boldsymbol{\mathfrak{s}}^{\pm}_{1}/\boldsymbol{\mathfrak{s}}^{\pm}_{0}$ determines a mapping class relation between products of positive Dehn twists on the Weinstein manifold
$(F,\omega_{F})=(Z^{+}_{1}\setminus B,\omega^{+}|_{TZ^{+}})$
as follows: Write $\operatorname{Critv}^{\pm}=\operatorname{Critv}\pi^{\pm}$ for the critical values of our holomorphic functions. We have a collection $\operatorname{\mathcal{L}}^{+}_{i}$ of Lagrangian spheres in $(F,\omega_{F})$ determined by the holomorphic function $\pi^{+}=\boldsymbol{\mathfrak{s}}^{+}_{1}/\boldsymbol{\mathfrak{s}}^{+}_{0}$ and a choice of matching paths for $\operatorname{Critv}^{+}$. We also have a collection $\phi^{-1}\operatorname{\mathcal{L}}^{-}_{i}$ in $(F,\omega_{F})$ determined by $\pi^{-}$ and a choice of matching paths for $\operatorname{Critv}^{-}$. Then appealing to Theorem 2.4.2, we have
$\tau_{L_{|\vec{L}^{+}|}^{+}}\cdots\tau_{L_{1}^{+}}=\tau_{\phi^{-1}L_{|\vec{L}^{-}|}^{-}}\cdots\tau_{\phi^{-}L_{1}^{-}}\in\pi_{0}\operatorname{Symp}_{0}(F,\omega_{F}).$
In other words, $\vec{L}^{+}$ and $\phi^{-1}\vec{L}^{-}$ gives positive relation. The difference in the lengths of these factorizations is easily counted from Theorem 2.4.2 as
$|\vec{L}^{+}|-|\vec{L}^{-}|=\#\operatorname{Crit}^{+}-\#\operatorname{Crit}^{-}=(-1)^{\dim_{\mathbb{C}}X^{\pm}}(\chi(X^{+})-\chi(X^{-})).$
The formula tells us that whenever we have a pencil pair with $\chi(X^{+})\neq\chi(X^{-})$, then their associated mapping class relations are non-braid-like.
3.2. The case $\dim_{\mathbb{C}}X^{\pm}=2$
Suppose that $X$ is a closed variety of $\dim_{\mathbb{C}}X=2$ with an ample line bundle $\operatorname{\mathcal{L}}$ having a pencil with associated holomorphic map $\pi$. Then our $Z_{1}$ is a Riemann surface and $B\subset Z_{1}$ is a set of points. Applying Poincaré duality and the adjunction formula,
#$\begin{gathered}\#B=[Z_{1}]^{2}=\int_{Z}\omega_{\operatorname{\mathcal{L}}}=\int_{X}c_{1}(\operatorname{\mathcal{L}})^{2},\quad\chi(Z_{1})=\int_{Z}c_{1}(X)-[Z_{1}]^{2}=\int_{X}(c_{1}(X)-c_{1}(\operatorname{\mathcal{L}}))\wedge c_{1}(\operatorname{\mathcal{L}})\\ \implies\#\operatorname{Crit}\pi=\int_{X}c_{2}(X)+(c_{1}(\operatorname{\mathcal{L}})-2c_{1}(X))\wedge c_{1}(\operatorname{\mathcal{L}}).\end{gathered}$
Since the symplectomorphism type of $(Z_{1},\omega_{\operatorname{\mathcal{L}}},B)$ is then determined by $\chi(Z_{1})$, the symplectic area of $Z_{1}$, and #$\#B$ (which is redundant), the following theorem is obvious.
Theorem 3.2.1.
Let $X^{\pm}$ be a pair of $\dim_{\mathbb{C}}=2$ varieties with ample line bundles $\operatorname{\mathcal{L}}^{\pm}$. Then there is a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair iff the quantities
$\int_{X^{\pm}}c_{1}(\operatorname{\mathcal{L}}^{\pm})^{2},\quad\int_{X^{\pm}}c_{1}(\operatorname{\mathcal{L}}^{\pm})c_{1}(X^{\pm})$
are independent of the $\pm$ signs.
Using the fact that $c_{1}(\operatorname{\mathcal{L}}^{\otimes}k)=kc_{1}(\operatorname{\mathcal{L}})$, we obtain the following easy lemma.
Theorem 3.2.2.
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ be a pencil pair with $\dim_{\mathbb{C}}X^{\pm}=2$. Then for all $k\geq 1$, there are $\phi_{k}$ for which the $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k},\phi_{k})$ are pencil pairs.
3.3. Generalized lantern relations
The following family of example is well-known: Consider $X^{+}=\mathbb{P}^{2},c_{1}(\operatorname{\mathcal{L}}^{+})=2\omega_{\mathbb{P}^{2}}$ and $X^{-}=\mathbb{P}^{1}\times\mathbb{P}^{1},c_{1}(\operatorname{\mathcal{L}}^{-})=(2\pi_{1}^{\ast}+\pi_{2}^{\ast})\omega_{\mathbb{P}^{1}})$. A quick calculation shows that these form a pencil pair by Theorem 3.2.1. The $Z^{\pm}_{1}$ are $\mathbb{P}^{1}$s each having self-intersection number $4$. The contact $3$-manifold given by the boundary of their tubular neighborhoods of the $Z^{\pm}_{1}\subset X^{\pm}$ is the standard lens space $(L(4,1),\xi_{std})$ given by reducing the standard contact $3$-sphere $(S^{3},\xi_{std})=\partial(\mathbb{D}^{4},\beta_{0})$ by the $\mathbb{Z}/4\mathbb{Z}$ action generated by $(z\in\mathbb{S}^{3})\mapsto iz$. The $X^{\pm}\setminus Z^{\pm}_{1}=W^{\pm}$ are its unique symplectic fillings (modulo symplectic blow-up) according to McDuff [M90]. The Lefschetz fibration on $W^{+}$ determined by a holomorphic Lefschetz pencil has fiber a $4$-punctured sphere with three critical points while the Lefschetz fibration on $W^{-}$ is a $4$-punctured sphere with four critical points. As observed by Auroux and Smith [AS08, §5.2] the Lefschetz fibration on $W^{+}$ has monodromy given the left hand side of the lantern relation as it appears in Figure 1. An analysis similar to that of [AS08] reveals that the monodromy of $W^{-}$ is given by the right-hand side of the lantern relation as it appears in Figure 1. From Theorem 3.2.2 we get pencil pairs from the $(X^{+},(\operatorname{\mathcal{L}}^{+})^{\otimes k})$ and $(X^{-},(\operatorname{\mathcal{L}}^{-})^{\otimes k})$ for all $k\in\mathbb{Z}_{>0}$.
Now we describe how the above $k\in\mathbb{Z}_{>0}$ family fits into a larger family of examples, starting with some generalities. Let $\operatorname{\mathcal{L}}_{\Sigma}$ be a line bundle of degree $-d=-d(\operatorname{\mathcal{L}}_{\Sigma})<-2g_{\Sigma}-1$ over a closed Riemann surface $\Sigma$. The $\operatorname{rank}=2$ bundle $\operatorname{Hom}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})=(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})^{\ast}\simeq\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma}^{\ast}$ over the total space of $\mathbb{C}\otimes\operatorname{\mathcal{L}}_{\Sigma}$ descends to a line bundle $\operatorname{\mathcal{L}}$ over the projectivization $\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})$ which is a ruled complex surface $\mathbb{P}^{1}\rightarrow\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})\rightarrow\Sigma$. Sections $\boldsymbol{\mathfrak{s}}$ of $\operatorname{\mathcal{L}}$ take the form $\boldsymbol{\mathfrak{s}}[z_{0},z_{1}]=az_{0}-\boldsymbol{\mathfrak{s}}_{\operatorname{\mathcal{L}}}(z_{1})$ where $\boldsymbol{\mathfrak{s}}_{\operatorname{\mathcal{L}}}$ is a section of $\operatorname{\mathcal{L}}_{\Sigma}^{\ast}\rightarrow\Sigma$, $z_{0}\in\mathbb{C}$, and $z_{1}\in\operatorname{\mathcal{L}}_{\Sigma}$. The condition on $-d$ implies that $\operatorname{\mathcal{L}}_{\Sigma}^{\ast}$ and so $\operatorname{\mathcal{L}}$ are ample. The zero locus of a $a=0$ section is the union of the section $[1,0]$ with $-d$ fibers of the fibration. Desingularizing such a non-generic section, the zero locus $Z$ of a generic section has the topological type of $\Sigma$. Since the section $[1,0]$ has self-intersection $-d$ and the fibers have self-intersection 00, $[Z]\bullet[Z]=d$. Since $\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma})$ is a $2$-sphere bundle with a section, $\chi(\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma}))=2\chi(\Sigma)$, eg. from a Gysin sequence. It follows that a pencil for $(\mathbb{P}(\mathbb{C}\oplus\operatorname{\mathcal{L}}_{\Sigma}),\operatorname{\mathcal{L}})$ has fiber $\Sigma$ minus $d$ points with $d$ critical points. It’s not difficult to work out that each singular fiber is a nodal curve given by a union of $\Sigma$ with $d-1$ points removed together with a fiber with one point removed. So we see that the total monodromy is a fibered Dehn twist on $\Sigma$ minus $d$ points. In general a pencil for $\operatorname{\mathcal{L}}^{\otimes k}$ the fiber $F_{k}$ is obtained by removing $k^{2}d$ punctures from a closed surface of $\chi=k(\chi(\Sigma)+d)-k^{2}d$. The Lefschetz fibration $\pi_{k}$ will have $2((1-k)\chi(\Sigma)-d)+3k^{2}d$ critical points.
Now consider $\Sigma_{l}$ a degree $l$ curve in $\mathbb{P}^{2}$ (so that $\chi(\Sigma)=l(3-l)$) with $\operatorname{\mathcal{L}}_{\Sigma}^{\ast}=\mathcal{O}_{k}|_{\Sigma_{k}}$ and consider $(X_{l},\operatorname{\mathcal{L}}_{l})$ as so that $d=l^{2}$. The fiber $F_{l,k}$ of a pencil for $(X_{l},\operatorname{\mathcal{L}}_{l}^{k})$ is then obtained from removing $(kl)^{2}$ points from a closed Riemann surface of $\chi=k(l(3-l)+l^{2})-(kl)^{2}$, with the associated Lefschetz fibration $\pi_{l,k}$ having #$\#\operatorname{Crit}\pi_{l,k}=2((1-k)l(3-l)-l^{2})+3(kl)^{2}$. In the case $l=2$, $(X_{l},\operatorname{\mathcal{L}}_{l})$ is the second Hirzebruch surface, being identical to the $(X^{-},\operatorname{\mathcal{L}}^{-})$ above. So for each such $l\geq 2$, the $(\mathbb{P}^{2},\mathcal{O}_{l})$ and $(X_{l},\operatorname{\mathcal{L}}_{l})$ determine a pencil pair. Applying Theorem 3.2.2, the $(\mathbb{P}^{2},\mathcal{O}_{l}^{\otimes k}=\mathcal{O}_{lk})$ and $(X_{l},\operatorname{\mathcal{L}}_{l}^{\otimes k})$ then determine pencil pairs for all $l\geq 2,k\geq 1$, summarized in the following theorem.
Theorem 3.3.1.
Let $F_{l,k}$ be an oriented surface of genus $g_{l,k}=1+\frac{1}{2}kl(kl-3)$ with $(kl)^{2}$ boundary components. Then a boundary Dehn twist $\tau_{\partial F_{l,k}}$ can be factorized as products of
  1. (1)
    $(kl)^{2}$ positive Dehn twists along the individual boundary components of $F_{k,l}$ determined by $\tau_{\partial F_{k,l}}$,
  2. (2)
    $3(kl)^{2}-6kl-2l^{2}+6l$ positive Dehn twists determined by a pencil for $(X_{l},\operatorname{\mathcal{L}}_{l}^{k})$, and
  3. (3)
    $3((kl)^{2}-3(kl)+1)$ positive Dehn twists determined by a pencil $(\mathbb{P}^{2},\mathcal{O}_{kl})$.
For the $k=1$ case the first two collections of curves coincide, and when $l=2,k=1$ these collections give the classical lantern relation.
Factorizing a given $kl\in\mathbb{Z}_{>1}$ in different ways then gives us many factorizations of $\tau_{\partial F_{l,k}}$. For example when $k=2^{n-j}$ and $l=2^{j}$ for $j=1,\dots,n-1<n$ the above theorem gives us factorizations into
$m=2^{n},3(2^{2n}-2^{n+1})+2^{j+1},3(2^{2n}-3\cdot 2^{n}+1)$
Dehn twists. By looking at the compliments of $\operatorname{\mathcal{L}}^{2^{n-j}}_{2^{j}}$ hyperplane sections in the $X_{l}$ and $\mathcal{O}_{2^{n}}$ hyperplane sections in $\mathbb{P}^{2}$, we see that the boundary of a tubular neighborhood of a $\deg=2^{n}$ curve in $\mathbb{P}^{2}$ has at least $n$ distinct (algebraic) Weinstein fillings, distinguishable by their Euler characteristics.
4. Pencil pairs from projective models
Our technique for finding new pencil pairs in this article will rely on our identification of the $Z^{\pm}_{1}\subset X^{\pm}$ with complete intersections in “big” projective varieties $\mathcal{M}$. In this section we describe this technique and use it show a infinite families of pencil pairs can be determined from a single pencil pair, analogous to the case of Theorem 3.2.2.
4.1. Projective models
Definition 4.1.1.
Let $(Z,\operatorname{\mathcal{L}})$ be a polarized variety having $\dim_{\mathbb{C}}Z=n$, let $\mathcal{M}\subset\mathbb{P}^{N}$ be a projective variety of $\dim_{\mathbb{C}}\mathcal{M}=m\geq n$, and let deg=($\deg=(d_{1},\dots,d_{m-n})\in\mathbb{Z}^{m-n}_{>0}$. We say that $(\mathcal{M},\deg)$ is a projective model for $(Z,\operatorname{\mathcal{L}})$ if there is a a holomorphic section $\boldsymbol{\mathfrak{s}}_{\mathcal{M}}$ of $\mathcal{O}_{\deg}=\oplus_{1}^{m-n}\mathcal{O}_{d_{j}}\rightarrow\mathbb{P}^{N}$ such that
  1. (1)
    there is a biholomorphism $\phi:Z\rightarrow Z_{\mathcal{M}}=\{\boldsymbol{\mathfrak{s}}_{\mathcal{M}}=0\}\cap\mathcal{M}$, and
  2. (2)
    $\phi^{\ast}\mathcal{O}_{\deg_{1}}=\operatorname{\mathcal{L}}\rightarrow Z$.
Lemma 4.1.2.
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ be a pair of ample polarized varieties. Suppose that there are holomorphic sections $\boldsymbol{\mathfrak{s}}^{\pm}_{1}$ of $\operatorname{\mathcal{L}}^{\pm}$ for which
  1. (1)
    the zero loci $Z_{1}^{\pm}=\{\boldsymbol{\mathfrak{s}}^{\pm}_{1}=0\}\subset X^{\pm}$ are transversely cut out, and
  2. (2)
    the $(Z^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have a common projective model $(\mathcal{M},\deg)$.
Then there are pencils for the ±$\operatorname{\mathcal{L}}^{\pm}\rightarrow X^{\pm}$ and a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
Proof.
By the definition of projective models there are sections $\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},1}$ of $\mathcal{O}_{\deg}$ with zero loci $Z_{\mathcal{M}}^{\pm}\subset\mathcal{M}$ and biholomorphisms $\phi^{\pm}:Z^{\pm}_{1}\rightarrow Z_{\mathcal{M}}^{\pm}$ for which ±$\operatorname{\mathcal{L}}^{\pm}|_{Z_{1}^{\pm}}=\phi^{\ast}\mathcal{O}_{\deg_{1}}$. Choose sections $\boldsymbol{\mathfrak{s}}^{\pm}_{0}$ of the $\operatorname{\mathcal{L}}^{\pm}$ so that the $(\boldsymbol{\mathfrak{s}}^{\pm}_{0},\boldsymbol{\mathfrak{s}}^{\pm}_{1})$ are pencils for the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ with base loci $B^{\pm}=\{\boldsymbol{\mathfrak{s}}^{\pm}_{0}=\boldsymbol{\mathfrak{s}}^{\pm}_{1}=0\}\subset Z^{\pm}_{1}$. By the existence of the projective models, there are sections $\boldsymbol{\mathfrak{s}}_{\mathcal{M},0}^{\pm}$ of $\mathcal{O}_{\deg_{1}}\rightarrow\mathbb{P}^{N}$ such that $\phi^{\pm}B^{\pm}\subset\mathcal{M}$ is cut out by the section $\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},1}\oplus\boldsymbol{\mathfrak{s}}_{\mathcal{M},0}^{\pm}$ of $\mathcal{O}_{\deg}\oplus\mathcal{O}_{1}\rightarrow\mathbb{P}^{N}$.
Let $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},1}\oplus\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},0}$ be a $T\in[-1,1]$ parameter family of sections of $\mathcal{O}_{\deg}\oplus\mathcal{O}_{1}\rightarrow\mathbb{P}^{m}$ for which $\boldsymbol{\mathfrak{s}}^{\pm 1}_{\mathcal{M}}=\boldsymbol{\mathfrak{s}}_{\mathcal{M}}^{\pm}$, $\boldsymbol{\mathfrak{s}}^{\pm 1}_{\mathcal{M},0}=\boldsymbol{\mathfrak{s}}_{\mathcal{M},0}^{\pm}$ and such that the zero loci of the $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},1}$ and $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M}}\oplus\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},0}$ are transversely cut out for all $T$. The transversality conditions can be satisfied using the fact that failure of transversality is a $\operatorname{codim}_{\mathbb{C}}\geq 1$ condition in the projective spaces of sections $\mathbb{P}(H^{0}(\mathcal{O}_{\deg}\oplus\mathcal{O}_{1}))$. A parametric Darboux-Moser-Weinstein argument gives us symplectomorphism
$\psi:(\phi^{+}Z^{\pm}_{1},\omega_{\mathcal{O}_{1}},\phi^{+}B^{+}=\{\boldsymbol{\mathfrak{s}}^{+}_{\mathcal{M}}\oplus\boldsymbol{\mathfrak{s}}^{+}_{\mathcal{M},0}=0\})\rightarrow(\phi^{-}Z^{\pm}_{1},\omega_{\mathcal{O}_{1}},\phi^{+}B^{-}=\{\boldsymbol{\mathfrak{s}}^{-}_{\mathcal{M}}\oplus\boldsymbol{\mathfrak{s}}^{-}_{\mathcal{M},0}=0\}).$
So $\phi=(\phi^{-})^{-1}\psi\phi^{+}$ a symplectomorphism $(Z^{+}_{1},\omega_{\operatorname{\mathcal{L}}^{+}},B^{+})\rightarrow(Z^{-}_{1},\omega_{\operatorname{\mathcal{L}}^{-}},B^{-})$. ∎
Definition 4.1.3.
In the above notation, $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is determined by the projective model $(\mathcal{M},\deg)$.
4.2. Cabling pencil pairs
The following theorem allows us to determine an infinite family of pencil pairs from a single pencil pair when it is determined by a projective model.
Lemma 4.2.1.
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi^{\pm})$ be determined by the projective model $(\mathcal{M},\deg=(d_{1},\dots,d_{m-n}))$. Then for all $k\in\mathbb{Z}_{\geq 1}$ there are $\phi_{k}$ for which there exists pencil pairs $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k},\phi_{k})$ determined by the projective model $(\mathcal{M},\deg_{k}=(kd_{1},d_{2},\dots,d_{m-n}))$.
Proof.
To start we study the local topology of divisors associated to powers of a line bundle. Let $Z\subset X$ be the zero locus of some transversely cut out section $\boldsymbol{\mathfrak{s}}$ of an ample $\operatorname{\mathcal{L}}\rightarrow X$. Consider sections $\boldsymbol{\mathfrak{s}}_{k}$ of $\operatorname{\mathcal{L}}^{\otimes k}$ the form $\boldsymbol{\mathfrak{s}}_{k}=\boldsymbol{\mathfrak{s}}^{\otimes k}+\mathfrak{t}$ with $\mathfrak{t}$ a $\mathcal{C}^{1}$ small section of $\operatorname{\mathcal{L}}^{\otimes k}$. By the smallness of $\mathfrak{t}$, the zero locus $Z_{k}$ of $\boldsymbol{\mathfrak{s}}_{k}$ is contained in an arbitrarily small tubular neighborhood of $Z$ in $X$. Writing $B_{k}=Z_{k}\cap Z=\{\mathfrak{t}=0\}\subset Z$, we see that $Z\setminus B_{k}$ is an affine variety since it is the complement of a hyperplane section in a projective variety. Hence $Z\setminus B_{k}$ is Stein. So by [Si76, Corollary 2] it has a neighborhood in $X$ whose projection onto $Z\setminus B_{k}$ is holomorphic. Possibly after multiplying $\mathfrak{t}$ by a small constant (which has no effect on $B_{k}$), we can assume that $Z_{k}\setminus B_{k}$ is contained in this neighborhood. The projection of this tubular neighborhood to $Z\setminus B_{k}$ restricted to $Z_{k}\setminus B_{k}$ gives a $k$-fold covering $\pi_{k}:Z_{k}\setminus B_{k}\rightarrow Z\setminus B_{k}$. By compactifying, we obtain a $k$-fold branched covering $\overline{\pi}_{k}:Z_{k}\rightarrow Z$ with branch locus $B_{k}$.
We claim that the isomorphism type of $(Z_{k},\operatorname{\mathcal{L}}^{\otimes k})$ depends only on $(Z,B_{k})$. The section $\mathfrak{t}$ trivializes $\operatorname{\mathcal{L}}^{\otimes k}$ over $Z_{k}\setminus B_{k}$ through which a holonomy morphism $\psi:\pi_{1}(Z\setminus Z_{k})\rightarrow\mathbb{S}^{1}$ is determined. Covering space theory tells us that the cover $\pi_{k}$ is smoothly classified by $\psi$ and so $\overline{\pi}_{k}$ is classified by $\psi$.
Since $\pi_{k}$ is holomorphic and a covering map, it is a local biholomorphism. Identifying $Z_{k}$ with the closure of $Z_{k}\setminus B_{k}$ in $X$, we see that the biholomorphism type of $Z_{k}$ is completely determined by the pair $(Z,B_{k})$. Observing that $B_{k}\subset Z_{k}$ is the zero locus $\boldsymbol{\mathfrak{s}}|_{Z_{k}}:Z_{k}\rightarrow\operatorname{\mathcal{L}}$ and applying the line-bundle/divisor correspondence, we see that the isomorphism type of $(Z_{k},\operatorname{\mathcal{L}})$ depends only on $(Z,B_{k})$. Hence the desired claim follows for $(Z_{k},\operatorname{\mathcal{L}}^{\otimes k})$.
It follows that is $(\mathcal{M},\deg)$ is a projective model for $(Z,\operatorname{\mathcal{L}})$, then $(\mathcal{M},\deg_{k})$ is a projective model for $(Z_{k},\operatorname{\mathcal{L}}^{\otimes k})$. Applying $\pm$ signs, we see that if $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ are determined by a projective model, then we can use Lemma 4.1.2 to find $\phi_{k}$ for which the $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k},\phi_{k})$ are pencil pairs determined by the projective models $(\mathcal{M},\deg_{k})$. ∎
5. $\dim_{\mathbb{C}}X^{\pm}=3$ examples
Now we work out some new examples of positive mapping class relations coming from $\dim_{\mathbb{C}}X^{\pm}=3$ pencils by applying the results of §4. First, we briefly review some vocabulary and known results.
The anticanonical bundle of a variety $X$ is defined $A_{X}=\wedge_{\mathbb{C}}^{\dim_{\mathbb{C}}X}TX$ for which $c_{1}(A_{X})=c_{1}(X)$. If $A_{X}$ is ample, $X$ is Fano and if $A_{X}$ is holomorphically trivializable, then $X$ is Calabi-Yau. If $X$ is Fano,
  1. (1)
    its index, $\operatorname{ind}_{X}$ is the maximal $k$ for which there is an ample $\operatorname{\mathcal{L}}\rightarrow X$ such that $A_{X}=\operatorname{\mathcal{L}}^{\otimes k}$, and
  2. (2)
    a generic section of $A_{X}$ cuts out an anticanonical divisor $Z$ (which is Calabi-Yau).
A simply connected $\dim_{\mathbb{C}}=2$ Calabi-Yau variety is a $K_{3}$ surface. According to Kodaira, the diffeomorphism type is uniquely determined. See [H16] for this and additional background. For an ample line bundle $\operatorname{\mathcal{L}}$ over a $K_{3}$ surface $Z$, the genus is define as
$g(Z,\operatorname{\mathcal{L}})=1+\frac{1}{2}\int_{Z}c_{1}(\operatorname{\mathcal{L}})^{2}$
and computes the genus of a hypersurface in $Z$ cut out by a section of $\operatorname{\mathcal{L}}$. It follows a Hodge theory computation that $\chi(Z)=24$. The adjunction formula applied to anticanonical divisors that for a $\dim_{\mathbb{C}}X=3$ Fano $\int_{X}c_{1}(X)c_{2}(X)=24$ as well, so that the only interesting Chern numbers of a Fano $X$ are $\int_{X}c_{1}(X)^{3}=\int_{X}A_{X}^{3}$ and $\chi(X)$. Combining these observations with some arithmetic of Chern classes, we obtain the following lemma.
Lemma 5.0.1.
If $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a pencil associated to a $\operatorname{\mathcal{L}}$ on a $\dim_{\mathbb{C}}X=3$ variety has
#$\#\operatorname{Crit}\pi=-\chi(X)+\int_{X}2c_{2}(X)c_{1}(\operatorname{\mathcal{L}})-3c_{1}(X)c_{1}(\operatorname{\mathcal{L}})^{2}+2c_{1}(\operatorname{\mathcal{L}})^{3}.$
So when $\operatorname{\mathcal{L}}=A_{X}^{\otimes k}$ on a Fano $X$,
#$\#\operatorname{Crit}\pi=-\chi(X)+48k+k^{2}(2k-3)\int_{X}A_{X}^{3}.$
5.1. Classification of polarized $K_{3}$ surfaces of low genus
We will be interested in $\dim_{\mathbb{C}}X=3$ Fano varieties for which we have a very ample $\operatorname{\mathcal{L}}\rightarrow X$ satisfying $\operatorname{\mathcal{L}}^{\otimes\operatorname{ind}_{X}}=A_{X}$. This implies that $\operatorname{\mathcal{L}}|_{Z}$ is very ample as well for $Z\subset X$ an anticanonical divisor. The following theorem reviews classification results of such polarized $K_{3}$ surfaces $(Z,\operatorname{\mathcal{L}})$ by Saint-Donat [SD74] (for $g(Z,\operatorname{\mathcal{L}})\leq 5$) and Mukai [M87] (for $6\leq g(Z,\operatorname{\mathcal{L}})\leq 10$).
Theorem 5.1.1.
In the above notation, suppose that $g=g(Z,\operatorname{\mathcal{L}})\leq 10$ where it may be computed
$g=1+\frac{1}{2}\int_{Z}c_{1}(\operatorname{\mathcal{L}})^{2}=1+\frac{1}{2}\operatorname{ind}_{X}\int_{X}c_{1}(\operatorname{\mathcal{L}})^{3}=1+\frac{1}{2}\operatorname{ind}_{X}^{-2}\int_{X}A_{X}^{3}.$
Then $(Z,\operatorname{\mathcal{L}})$ has a projective model $(\mathcal{M}_{g}\subset\mathbb{P}^{N_{g}},\deg_{g})$ which is entirely determined by $g$.
Note that we require $\operatorname{\mathcal{L}}$ to be very ample, rather than just ample which is insufficient for the desired result [M87]. For $g=3$, $Z$ is a $\deg=4$ hypersurface in $\mathbb{P}^{3}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}$ so the model is $(\mathcal{M}_{3}=\mathbb{P}^{4}\subset\mathbb{P}^{4},\deg_{4}=(1,4))$. For $g=4$, $Z$ is an intersection of $\deg=2$ and $\deg=3$ hypersurfaces in $\mathbb{P}^{4}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}\rightarrow\mathbb{P}^{4}$, so the model is $(\mathcal{M}_{5}=\mathbb{P}^{5},\deg_{5}=(1,2,3))$. For $g=5$, $Z$ is an intersection of three $\deg=2$ hypersurfaces in $\mathbb{P}^{5}$ with $\operatorname{\mathcal{L}}=\mathcal{O}_{1}\rightarrow\mathbb{P}^{5}$, so $(\mathcal{M}_{5}=\mathbb{P}^{5},\deg_{5}=(1,2,2,2))$. In the $6\leq g\leq 10$ cases, the $\mathcal{M}_{g}$ have positive codimension in $\mathbb{P}^{N_{g}}$ and will be described in the following subsection.
5.2. Examples of pencil pairs from anticanonical divisors on Fano $3$-folds
Combining the above results with Lemmas 4.1.2 and 4.2.1, we have the following theorem.
Theorem 5.2.1.
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ be polarized Fano $3$-folds for which $A_{X^{\pm}}=(\operatorname{\mathcal{L}}^{\pm})^{\operatorname{ind}_{X^{\pm}}}$ and $\operatorname{\mathcal{L}}^{\pm}$ are very ample. Suppose that their anticanonical divisors $Z^{\pm}\subset X^{\pm}$ are such that $g(Z^{\pm},\operatorname{\mathcal{L}}^{\pm})\leq 10$, or equivalently
$\int_{X^{\pm}}A_{X^{\pm}}^{3}\leq 18\operatorname{ind}_{X^{\pm}}^{2}$
Then for every $k\in\mathbb{Z}_{>0}$ there are $\phi_{k}$ for which $(X^{\pm},A_{X^{\pm}}^{\otimes k},\phi_{k})$ are pencil pairs.
To find all examples of pencil pairs determinable from Theorem 5.2.1, we cross reference the above theorems with the complete classification of $\dim_{\mathbb{C}}=3$ Fano varieties by Iskovskih [I77, I77] and Mori-Mukai [MM82]. There are $105$ such varieties which are listed in the Fanography database [Be24] along with relevant data such as indices and Hodge diamonds (which of course determine Euler characteristics). [^1] See https://www.fanography.info/about for a complete list of resources from which the database is compiled.
We collect the relevant examples of our Fano $X$ into families (having at least two members) organized by their $\operatorname{ind}_{X}$ and $\int_{X}A_{X}^{3}$. Members of the families are labeled by their indices $a$-$b$ as in [Be24], with $a$ being the Picard rank. For each, a short description of $X$ is given. According to Theorem 5.2.1 any two members of a given family form a pencil pair.
5.2.1. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=10\implies g(Z,A_{X})=6$
  1. (1-5)
    $X$ is an intersection of the $\mathcal{M}_{6}=G(2,5)\subset\mathbb{P}^{9}$ by two hyperplanes and a quadric having $\chi=-16$.
  2. (2-4)
    $X$ is a divisor of $\deg=(1,3)$ on $\mathbb{P}^{1}\times\mathbb{P}^{3}$ having $\chi=-14$.
Here is the projective model is $(\mathcal{M}_{6},\deg=(1,1,1,2))$ following [M87].
We obtain Weinstein $4$-manifolds $(F_{k},\beta_{k})$ given by a $A_{X}^{\otimes k}$ divisor in either $X$, minus its intersection with another such generic divisor. Applying Theorem 5.2.1, we have factorizations of $\tau_{\partial F_{k}}$ into products of
$\chi+48k+10k^{2}(2k-3),\quad\chi=-16,-14$
Dehn twists along exact Lagrangian spheres.
5.2.2. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=12\implies g(Z,A_{X})=7$
  1. (1-6)
    $X$ is an intersection of a $\mathcal{M}_{7}\subset\mathbb{P}^{15}$ of $\dim_{\mathbb{C}}\mathcal{M}=10$ with four hyperplanes having $\chi(X)=-10$ as described in [M87].
  2. (2-5)
    $X$ is an intersection of $\deg=(0,3)$ and $\deg=(1,1)$ hypersurfaces in $\mathbb{P}^{1}\times\mathbb{P}^{4}$ having $\chi(X)=-6$.
  3. (2-6)
    $X$ is a $\deg=(2,2)$ divisor in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ having $\chi(X)=-12$.
  4. (3-1)
    $X$ is a double branched cover of $(\mathbb{P}^{1})^{3}$ over a $\deg=(2,2,2)$ divisor having $\chi(X)=-8$.
In this case the projective model of [M87] is $(\mathcal{M}_{7},(1,1,1,1,1))$.
Theorem 5.2.1 gives us Weinstein $4$-manifolds $(F_{k},\beta_{k})$ and factorizations of $\tau_{\partial F_{k}}$ into products of
$-\chi+48k+12k^{2}(2k-3),\quad\chi=-10,-6,-12,-8$
Dehn twists along exact Lagrangian spheres.
5.2.3. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=14\implies g(Z,A_{X})=8$
  1. (1-7)
    $X$ is a the intersection of $\mathcal{M}_{8}=G(2,6)\subset\mathbb{P}^{14}$ with nine hyperplanes having $\chi(X)=-6$.
  2. (2-7)
    $X$ is an intersection of $\deg=(0,2)$ and $\deg=(1,2)$ divisors on $\mathbb{P}^{1}\times\mathbb{P}^{4}$ having $\chi(X)=-4$.
  3. (2-8)
    $X$ is a double branched cover of $\mathbb{P}(\mathbb{C}\oplus\mathcal{O}_{1}\rightarrow\mathbb{P}^{2})$ over an anticanonical divisor with $\chi(X)=-12$.
  4. (3-2)
    $X$ is a divisor in a $\mathbb{P}^{2}$ bundle over $\mathbb{P}^{2}$ having $\chi(X)=2$.
Theorem 5.2.1 gives us Weinstein $4$-manifolds $(F_{k},\beta_{k})$ and factorizations of $\tau_{\partial F_{k}}$ into products of
$-\chi+48k+14k^{2}(2k-3),\quad\chi=-6,-4,-12,2$
Dehn twists along exact Lagrangian spheres.
5.2.4. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=16\implies g(Z,A_{X})=9$
  1. (1-8)
    $X$ is an intersection of a $\dim_{\mathbb{C}}=6$ spin Grassmann variety in $\mathbb{P}^{13}$ with three hyperplanes, having $\chi(X)=-2$.
  2. (2-9)
    $X$ is an intersection of $\deg=(1,1)$ and $\deg=(1,2)$ divisors in $\mathbb{P}^{2}\times\mathbb{P}^{3}$ having $\chi(X)=-4$.
  3. (2-10)
    $X$ is an intersection of $\deg=(0,2)$, $\deg=(0,2)$, and $\deg=(1,1)$ divisors in $\mathbb{P}^{1}\times\mathbb{P}^{5}$ having $\chi(X)=0$.
Again the first example corresponds to the $g=9$ model of [M87].
Theorem 5.2.1 gives us Weinstein $4$-manifolds $(F_{k},\beta_{k})$ and factorizations of $\tau_{\partial F_{k}}$ into products of
$-\chi+48k+16k^{2}(2k-3),\quad\chi=-2,-4,0$
Dehn twists along exact Lagrangian spheres.
5.2.5. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=18\implies g(Z,A_{X})=10$
  1. (1-9)
    $X$ is a $\dim_{\mathbb{C}}=5$ Grassmann bundle for the exceptional Lie group $G_{2}$ inside of $\mathbb{P}^{13}$, intersected with two hyperplanes having $\chi(X)=0$.
  2. (2-11)
    $X$ is a blowup of a $\deg=3$ hypersurface in $\mathbb{P}^{4}$ along a line having $\chi(X)=-4$.
  3. (3-3)
    $X$ is a $\deg=(1,1,2)$ divisor on $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{2}$ having $\chi(X)=2$.
  4. (3-4)
    $X$ is obtained by taking a branched cover of $\mathbb{P}^{1}\times\mathbb{P}^{2}$ along a $\deg=(2,2)$ divisor and blowing up along a smooth fiber of the projection to $\mathbb{P}^{2}$, having $\chi(X)=4$.
  5. (8-1)
    $X$ is a blowup of $\mathbb{P}^{2}$ at six points times $\mathbb{P}^{1}$ having $\chi(X)=18$.
Again, the first member of the family corresponds to the model of [M87].
Theorem 5.2.1 gives us Weinstein $4$-manifolds $(F_{k},\beta_{k})$ and factorizations of $\tau_{\partial F_{k}}$ into products of
$-\chi+48k+12k^{2}(2k-3),\quad\chi=-4,0,4,2,18$
Dehn twists along exact Lagrangian spheres. The case $k=1$ is Theorem 1.0.1.
5.2.6. Case $\operatorname{ind}_{X}=2,\int_{X}A_{X}^{3}=48\implies g(Z,\operatorname{\mathcal{L}})=7$
  1. (2-32)
    $X$ is a divisor of $\deg=(1,1)$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ with $\chi(X)=6$. Here $\operatorname{\mathcal{L}}$ is of degree $(1,1)$ so that $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Since $\operatorname{\mathcal{L}}$ is the pullback of $\mathcal{O}_{1}$ from the Segre embedding $\mathbb{P}^{2}\times\mathbb{P}^{2}\rightarrow\mathbb{P}^{8}$ and so is very ample.
  2. (3-27)
    $X=(\mathbb{P}^{1})^{3}$ having $\chi(X)=8$. Here $\operatorname{\mathcal{L}}$ has $\deg=(1,1,1)$ giving $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Again $\operatorname{\mathcal{L}}$ is very ample as it is determined by a Segre embedding.
These are exactly the Fano $3$-folds considered in [O21], where divisors cut out by $\operatorname{\mathcal{L}}$ – rather than $\operatorname{\mathcal{L}}^{\otimes 2}$ – are considered to create pencil pairs. Here Theorem 5.2.1 gives us Weinstein $4$-manifolds $(F_{k},\beta_{k})$ and factorizations of $\tau_{\partial F_{k}}$ into products of $-\chi+48(2k^{3}-3k^{2}+k),\chi=6,8$ Dehn twists along exact Lagrangian spheres. We will generalize these examples in the next section.
5.3. Oba’s example revisited
As above consider $(X^{+},\operatorname{\mathcal{L}}^{+})$ given by a $\deg=(1,1)$ hypersurface in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ equipped with $\operatorname{\mathcal{L}}^{+}$ a $\deg=(1,1)$ line bundle and $(X^{-},\operatorname{\mathcal{L}}^{-})$ is $(\mathbb{P}^{1})^{3}$ with $\operatorname{\mathcal{L}}^{-}$ a $\deg=(1,1,1)$ line bundle. Then the $(Z^{\pm}_{1},\operatorname{\mathcal{L}}^{\pm})$ are $\int A_{Z}^{2}=6$ del Pezzo surfaces (that is, Fano $2$-folds) with $\operatorname{\mathcal{L}}^{\pm}$ the anticanonical bundle $A_{Z^{\pm}_{1}}$. So smoothly $Z^{\pm}_{1}=\mathbb{P}^{2}\#3\overline{\mathbb{P}^{2}}=\left(\mathbb{P}^{1}\times\mathbb{P}^{1}\right)\#2\overline{\mathbb{P}^{2}}$.
Polarized del Pezzo surfaces $(Z,\operatorname{\mathcal{L}}=A_{X})$ of $\int A_{Z}^{2}=6$ are isomorphic by [F80]. Any such biholomorphism $\phi:Z^{+}_{1}\rightarrow Z^{-}_{1}$ lifts to an isomophism of anticanonical bundles, by the fact that the anticanonical bundles are constructed from the $TZ^{\pm}_{1}$. We can then consider either $(Z^{\pm}_{1},\operatorname{\mathcal{L}}^{\pm})$ as being a projective model and obtain a pencil pair $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ from Lemma 4.1.2. Applying $l$-fold cables as in 4.2.1, we get pencil pairs $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes l},\phi_{l})$ for all $l\geq 1$ which recover Oba’s example [O21] in the case $l=1$, and recover the examples of §5.2.6 when $k=2l$. So the $l=2$ case gives us the polarized $K_{3}$ surfaces $(Z^{\pm}_{1},(\operatorname{\mathcal{L}}^{\pm})^{2}=A_{X})$ of $g=7$.
For the associated mapping class relations, we have $(F_{l},\beta_{l}$ given by a $\deg=(l,l,l)$ hypersurface in $(\mathbb{P}^{1})^{3}$ minus its intersection with another hypersurface of the same degree. A quick computation counts the number of Dehn twists of the associated mapping class relation as
$812l(2l^{2}-3l+2)-\chi(X^{\pm}),\quad\chi(X^{+})=6,\quad\chi(X^{-})=$ 8 .